14. Kleinfeldt Company allocates overhead on the basis of direct labor hours. It allocates overhead costs of $4,000 to two different jobs as follows: Job 1: (10 hours)- $2,000 Job 2: (10 hours) $2,000 The production process for Job 1 is automated. Now Job 1 requires only 2 hours of direct labor but four hours of mechanical processing. As a result, total overhead increased to $6,000. How much overhead cost will be assigned to Job 2 after automation? A) $3,000 B) $5,000 C) $2,000 D) $1,000 15. Aaron Company makes two products. Making Product X requires 4,000 hours of labor, and Product Y requires 6,000 hours of labor. Aaron undertook an automation program that reduced the consumption of labor required by Product Y to only 2,000 hours of labor. Product X was not affected by the automation process. Overhead cost prior to the automation totaled $10,000. After automation, overhead cost amounted to $24,000. Assuming Aaron uses direct labor hours as a company-wide allocation base before and after the automation, the amount of overhead cost allocated to A) Product X would be $4,000 prior to automation and $16,000 after automation. B) Product X would be $6,000 prior to automation and $8,000 after automation. C) Product Y would be $8,000 prior to automation and $8,000 after automation. D) Product Y would be $2,000 prior to automation and $12,000 after automation. 16. Setup costs for XYZs two products amounted to $2,000. The company used 2,000 direct labor hours to produce product A and 1,200 direct labor hours to produce product B. Both products required the same number of setups. XYZ uses a company-wide overhead rate to allocate overhead based on direct labor hours. Based on this information A) Product A is undercosted by $250 under traditional costing. B) Product A is overcosted by $250 under traditional costing. C) Product B is undercosted by $750 under traditional costing. D) Product B is overcosted by $250 under traditional costing.
Answer #1
Requirement 14 :
Particulars Job 1 Job 2 Total labour Hour
2 10 12
overhead  allocated based on Direct labor hour 6000/12 = 500 per labour hour
JOB 1 JOB 2
2 10
Overhead allocated @ 500 per labour hour 1000 5000
Option [c] is correct.
Requirement 15 :
Option[a] is correct.
Assuming aaron uses direct labour hours as a company wide allocation base before and after the automation , The amount of overhead allocated is :
Overhead costs prior to automation $10,000 X=4000 hours Y=6000 hours
X=$4000 Y=$6000
Overhead costs after to automation $24,000 X=4000 hours Y=2000 hours
X=24000/6000*4000 Y=24000/6000*2000
$ 16,000.00 $ 8,000.00
Product X would be $4000 prior to automation and $ 16000 after automation
Requirement 16 :
Set up costs = $2000
Total number of direct labour hours = 3200 hours
Company wide overhead rate = 2000/3200 0.625 Per labour hour
Hours Rate per labour hour
Product A 2000 0.625 1250
Product B 1200 0.625 750
If number of set ups are same, the set up cost should be $1000 each.
Therefore,Product A is overcosted by $250 under traditional costing.
Option[b] is correct.